A radiator contains 10 quarts of fluid, 30% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 60% antifreeze?
Answer by stanbon(75887)

A radiator contains 10 quarts of fluid, 30% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 60% antifreeze?
----
Equation:
anti - anti + anti = anti
------
0.3*10 - 0.3*x + 1*x = 0.6*10
0.7x = 0.3*10
x = (3/7)10
x = 4.29 quarts (amt. to remove and replace)
----------
Cheers,
Stan H.